OpenStudy (anonymous):

Use the substitution method to solve the system of equations. Enter your answer as an ordered pair. 3x - y = 7 2x - 2y = 2

OpenStudy (anonymous):

i would multiply the top equation by -2 and then add them together to see what you get

OpenStudy (anonymous):

3x-y=7 -y=7-3x y=-7+3x sub to 2x-2y=2 2x-2(-7+3x)=2 2x-6x+14=2 -4x=-12 x=3 sub to y=-3x-7 y=-3(3)-7 y=-16 (3,-16)

OpenStudy (anonymous):

2x-2y = 2 implies x-y = 1 implies x = y+1 3x-y = 7 implies 3(y+1)-y=7 implies 3y+3-y=7 im;ies 2y=4 implies y=2 x=2+1 =3 x,y = 3,2

OpenStudy (anonymous):

(3,2)

OpenStudy (anonymous):

multiply first equation throughout by 2

OpenStudy (anonymous):

6x-2y=14 -)2x-2y=2 -------- 4x=12

OpenStudy (anonymous):

x=3 no wsub that into original eqn

OpenStudy (anonymous):

u'll get 2

OpenStudy (anonymous):

i correct myself 3x-y=7 -y=7-3x y=-7+3x sub to 2x-2y=2 2x-2(-7+3x)=2 2x-6x+14=2 -4x=-12 x=3 sub to y=-7+3x y=-7+9 y=2 (3,2)